# See What You Seek: Semantic Contextual Integration for Cloth-Changing Person Re-Identification

### v1

@@ Abstract
Cloth-changing person re-identification (CC-ReID) aims to match individuals across multiple surveillance cameras despite variations in clothing. Existing methods typically focus on mitigating the effects of clothing changes or enhancing ID-relevant features but often struggle to capture complex semantic information. In this paper, we propose a novel prompt learning framework, Semantic Contextual Integration (SCI), for CC-ReID, which leverages the visual-text representation capabilities of CLIP to minimize the impact of clothing changes and enhance ID-relevant features. Specifically, we introduce Semantic Separation Enhancement (SSE) module, which uses dual learnable text tokens to separately capture confounding and clothing-related semantic information, effectively isolating ID-relevant features from distracting clothing semantics. Additionally, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the model’s focus on distinctive ID characteristics. This integration enhances the model’s discriminative power and enriches the visual context with high-dimensional semantic insights. Extensive experiments on three CC-ReID datasets demonstrate that our method outperforms state-of-the-art techniques. The code will be released at github.
@@
Person re-identification (Re-ID) is a critical task that involves identifying individuals across different locations over time, with significant applications in video surveillance and smart city infrastructure [1, 2, 3]. Traditional Re-ID methods [4, 5, 6, 7] focus on clothing features, such as texture and color, to improve performance. However, these methods become ineffective when individuals change clothes, highlighting the need for robust approaches that can handle clothing variations in real-world settings.
@@
To address the challenge of clothing changes in Re-ID (CC-ReID), extensive research has yielded promising results [8, 9]. Existing methods can be broadly categorized into two categories: those aiming to reduce the influence of clothing and those focusing on enhancing ID-relevant information. The first category seeks to minimize the impact of clothing features, as illustrated in Fig. 1(a). For instance, CAL [10] uses adversarial learning to reduce reliance on clothing by penalizing predictions based on the RGB modality, while AIM [11] employs causal intervention with a dual-branch model to mitigate clothing bias. The second category emphasizes ID-relevant features using auxiliary cues such as pose, gait, or human parsing networks. For example, GI-ReID [12] uses gait information to learn clothing-agnostic representations, and SCNet [13] leverages human body parsing network to learn identity-relevant features through semantic consistency constraints.
@@
Although these methods address clothing changes, they typically focus either on minimizing clothing influence or enhancing ID cues through explicit features like body contours. We argue that implicit positive factors such as hair, glasses, or backpacks are embedded in the visual features but are challenging to define explicitly or separately. As shown in Fig. 1(b), our goal is to remove negative influences while making implicit positive factors more explicit.
Recent advancements in visual-language learning have shown great potential in enhancing visual recognition and understanding. Contrastive Language-Image Pre-training (CLIP) [14], a prominent cross-modal pre-training model, has effectively bridged the gap between visual data and natural language, enabling more context-aware and interpretatively rich analysis. CLIP has achieved remarkable success in various downstream visual tasks [15, 16]. In Re-ID, recent work [17, 18] demonstrates that linking visual content with corresponding language descriptions allows models to capture complex semantics associated with individuals. However, these studies have primarily focused on traditional Re-ID, applying CLIP to CC-ReID has not yet been explored. This motivates us to leverage prompt learning to reveal high-dimensional semantic information, thereby enhancing visual-textual connections and making implicit positive factors more explicit.
In this paper, we propose the Semantic Contextual Integration (SCI) framework for CC-ReID, which utilizes CLIP to capture high-dimensional semantic features. Since CLIP struggles to selectively ignore aspects like clothing through negative prompts, we introduce a Semantic Separation Enhancement (SSE) module to minimize the influence of clothing by isolating semantic-level negative factors and preserving key positive features for CC-ReID.
Specifically, we use dual learnable text tokens to separately capture confounding and clothing-related semantic information. We then apply an orthogonalization process on two distinct sets of text features to filter out clothing-related features from the confounding ones. SSE not only reduces the negative impact of clothing but also retains important positive semantic factors. These refined text features are used to guide the visual encoder.
To further emphasize implicit positive factors in visual images, we introduce a Semantic-Guided Interaction Module (SIM), which directs the visual encoder’s extraction process using the filtered text features. This allows the model to focus on ID-relevant elements beyond explicit features such as body contours. By leveraging CLIP’s text-visual integration capabilities, our method enhances the interaction between the two modalities, improving visual feature refinement and multi-modal alignment.
Our contributions are summarized fourfold:
@@
Extensive experiments on three standard CC-ReID datasets demonstrate the effectiveness of our approach, providing a robust solution to real-world Re-ID challenges.
@@ II Related Work
II-A Cloth-Consistent Person Re-ID
Person Re-ID under consistent clothing conditions has been extensively studied [19, 20, 21]. These methods primarily rely on clothing color and texture to extract discriminative features. Feature representation learning in this context can be categorized into three types: global features, local features, and auxiliary information-based features.
Global feature representation methods extract a single feature vector for each person image [22]. Local feature representation methods aggregate part-based features to address misalignments, using techniques like human parsing [23] for semantic part detection, horizontal division [24] for structured feature aggregation, hard region-level and soft pixel-level attention [25] for optimized features, and IANet [26] for enhanced small visual cues. Auxiliary information-based methods incorporate additional data, such as semantic attributes [19] or generated/augmented samples [20], to enhance generalization by embedding richer contextual information. And ISP [27] locates human body parts and personal belongings at pixel-level for aligned with labels.
Although effective in handling variations in appearance, background, and camera perspectives, these methods assume consistent clothing, leading to performance declines in scenarios with clothing changes, which are common in real-world applications. This limitation has spurred increased focus on addressing cloth-changing person Re-ID challenges.
II-B Cloth-Changing Person Re-ID
To advance CC-ReID, several datasets such as PRCC [28], LTCC [29], Celeb-reID [30], and VC-Clothes [31] have been introduced, providing diverse data for effective model evaluation. The core challenge in CC-ReID is to learn features that reliably identify individuals despite clothing variations.
There are three primary approaches to this challenge. The first leverages auxiliary soft-biometric information. For example, CAMC [32] incorporates body shape semantics into ID-related features, enhancing robustness to clothing changes, while FSAM [33] learns fine-grained body shape features to complement clothing-independent features. M2Net [34] utilizes contour and human parsing information to learn features that are robust to appearance changes.
The second approach involves disentangling features to isolate clothing-irrelevant information and reduce clothing bias. For instance, CAL [10] uses clothes-based adversarial loss to emphasize clothing-irrelevant features, and AIM [11] applies causal intervention to automatically eliminate clothing effects.
The third approach employs data augmentation or adaptation techniques. Pos-Neg [8] and CCFA [9] enhance clothing color and texture diversity without additional data collection, improving robustness to clothing variations. RCSANet [35] explicitly constructs a clothing status, which is used to enhance the robustness of ID features for handling both cloth-changing and no-cloth-changing cases.
Despite these advancements, existing methods primarily rely on visual inputs for model training, limiting their capacity for deep semantic understanding. In this paper, we propose a novel method that integrates visual and natural language information, utilizing CLIP to enhance the interaction of the visual content with corresponding textual descriptions.
@@ II-C Vision-Language Learning
Vision-language pre-training (VLP) has recently shown significant promise in enhancing the performance of various downstream vision tasks by training models to associate images with textual information. CLIP [14], a leading VLP model, employs paired image and text encoders to bridge the gap between natural language and visual content. By using similarity learning, CLIP aligns textual and visual representations in a shared space, enhancing tasks such as image captioning and classification.
Building on CLIP, recent advances in prompt learning, such as CoOp [36], transform prompt context words into learnable vectors. CoCoOp [37] extends this by generating input-conditional tokens for each image through a lightweight network. DenseCLIP [38] adapts this approach for dense prediction models, introducing pixel-text matching to utilize contextual information from a pre-trained language model.
In person Re-ID, CLIP-ReID [17] is the first to use ID-specific learnable tokens and a two-stage training strategy to fully utilize text for specifying visual concepts. CCLNet [18] advances this approach with a novel prompt learning paradigm for unsupervised visible-infrared Re-ID. However, these studies primarily focus on scenarios with consistent clothing, leaving CLIP’s potential in CC-ReID largely unexplored. In this paper, we propose the CLIP-CCReID framework to mitigate the impact of clothing changes in CC-ReID.
@@ III-A Preliminaries and Overview
The large-scale pre-trained vision-language model CLIP [14] demonstrates significant advantages in learning joint representations of images and text. CLIP consists of two encoders: a visual encoder and a text encoder . CLIP aligns the embedding spaces of visual and textual data during pre-training using a contrastive objective.
For downstream classification tasks, a simple and effective method [36] constructs text prompts using a template like “a photo of a [CLS].”, where [CLS] is replaced with specific class names. Given an image, CLIP computes similarities between the image embedding and the text prompt embeddings, the process involves transforming the image and text into a shared high-dimensional space where their relationships can be quantitatively assessed. CLIP then selects the class corresponding to the highest similarity as the prediction.
Building upon CLIP, CoOp [36] introduces learnable textual contexts, represented as “ [CLS].”, where () are learnable tokens. This approach enhances the model’s transferability and performance. Leveraging the capabilities of CLIP and CoOp, we employ these textual descriptions to achieve deeper semantic understanding for CC-ReID, as illustrated in Fig. 2, enabling more robust and accurate modeling despite clothing changes.
@@ III-B Semantic Separation Enhancement
Since CLIP cannot learn negative semantics or ignore specific elements like clothing through negative prompts, we propose the Semantic Separation Enhancement (SSE) module to address this limitation. This module mitigates the impact of clothing variations by subtracting negative semantic elements, preserving only the high-dimensional positive features crucial for CC-ReID. The complete algorithm is summarized in Algorithm 1.
We define dual textual prompts as follows:
@@
We then perform orthogonalization by subtracting the projection from , removing the component in the direction of at the textual level. This process ensures that the features are aligned with the positive features while removing the influence of negative factors. The orthogonalization process is:
where is the feature orthogonalized to align with and orthogonal to .
@@ III-C Semantic-Guide Interaction Module
Previous work [10, 17, 39] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios. However, these methods typically treat visual and textual branches independently, lacking interaction. Our method refines visual representations by leveraging clothing-irrelevant textual features, enhancing the interaction between visual and textual branches. This integration introduces more robust and invariant descriptors into the visual processing, improving the model’s ability to identify individuals across clothing changes. The algorithm is summarized in Algorithm 2.
As illustrated in Fig. 3, we employ a cross-attention mechanism in a Transformer decoder [40] to model interactions between visual and textual data. Instead of traditional self-attention, we use a non-local operation [41] to capture global context, ensuring the model can integrate information from distant but relevant positive parts of the visual image, which is crucial in cloth-changing scenarios. During this stage, we update the visual encoder while keeping other components frozen. The computation is as follows:
@@
We further enhance interaction by using text embeddings as queries and visual embeddings as keys and values. This cross-attention mechanism refines visual features using text information, ensuring alignment with relevant semantic context:
where is the dimensionality of the keys. This implementation strategically encourages the text features to find the most relevant visual clues.
@@
This module employs an interaction-to-alignment approach, ensuring that text embeddings guide the visual encoding process, transforming implicit information into explicit visual cues. By emphasizing positive factors and reducing negative ones, this method enhances the model’s ability to consistently identify individuals across various clothing scenarios.
@@
where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , and is a temperature parameter.
@@ IV-A Datasets and Evaluation Metrics
We evaluate our proposed method on three standard cloth-changing datasets: LTCC [29], PRCC [28], and VC-Clothes [31], following CAL [10] and AIM [11]. Additionally, we also evaluate our method on two traditional Re-ID datasets: Market-1501 [42], and MSMT17 [43], to validate its applicability.
LTCC dataset contains 17,138 images of 152 IDs captured by 12 cameras. The training set comprises 9,576 images of 77 IDs, and the testing set includes 7,543 images (493 queries and 7,050 gallery images) of 75 IDs. This long-term Re-ID dataset features frequent clothing changes and varied environmental conditions, with each ID having between 2 and 14 outfits.
PRCC dataset consists of 33,698 images of 221 IDs captured by three cameras. The training set contains 22,898 images of 150 IDs, while the testing set includes 10,800 images of 71 IDs. Each person has two outfits, with cameras A and B capturing the same clothes and camera C capturing a different outfit.
VC-Clothes dataset is a synthetic dataset generated using the GTA5 game engine. It comprises 9,449 images of 256 IDs across both training and testing sets, featuring 1,241 clothing items across four camera views.
Market-1501 dataset includes 1,501 individuals recorded by six different cameras. It consists of 12,936 images of 751 IDs for training and 19,732 images of 750 IDs for testing. The evaluation is conducted using the single query setting.
MSMT17 dataset includes 1,041 individuals captured by 15 different cameras. The training set contains 32,621 images of 1,041 IDs, while the testing set includes 51,027 images of 11,659 IDs. This dataset is notably large and complex.
We use the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics. Rank- in CMC measures the likelihood of retrieving the correct ID in the top- results, while mAP captures the average retrieval performance across all queries.
Evaluation is conducted under three settings: 1) General setting: Excludes same-ID and same-camera samples, using both cloth-changing and clothing-consistent samples. 2) Same-clothes setting: Excludes same-ID and same-camera samples, using only clothing-consistent samples. 3) Cloth-changing setting: Excludes same-ID, same-camera, and same-clothing samples, using only cloth-changing samples.

### v2

@@ Abstract
Cloth-changing person re-identification (CC-ReID) aims to match individuals across surveillance cameras despite variations in clothing. Existing methods typically mitigate the impact of clothing changes or enhance identity (ID)-relevant features, but they often struggle to capture complex semantic information. In this paper, we propose a novel prompt learning framework Semantic Contextual Integration (SCI), which leverages the visual-textual representation capabilities of CLIP to reduce clothing-induced discrepancies and strengthen ID cues. Specifically, we introduce the Semantic Separation Enhancement (SSE) module, which employs dual learnable text tokens to disentangle clothing-related semantics from confounding factors, thereby isolating ID-relevant features. Furthermore, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the focus of the model on distinctive ID characteristics. This semantic integration improves the discriminative power of the model and enriches the visual context with high-dimensional insights. Extensive experiments on three CC-ReID datasets demonstrate that our method outperforms state-of-the-art techniques. The code will be released at https://github.com/hxy-499/CCREID-SCI.
@@
Person re-identification (ReID) is the task of matching individuals across non-overlapping cameras over time, with important applications in video surveillance and smart city systems [1, 2, 3, 4]. Traditional ReID methods [5, 6] exploit the appearance of clothing, e.g., texture and color, to distinguish identities (IDs), but their performance degrades when subjects change clothing. This limitation motivates the development of cloth-changing ReID (CC-ReID) techniques that remain robust under clothing variations.
@@
Existing CC-ReID approaches can be grouped into two categories: those that suppress clothing cues and those that enhance ID-relevant features (see Fig. 1(a)). The first category seeks to minimize the influence of clothing: CAL [7] employs adversarial learning to penalize clothing-based predictions in the RGB modality, and AIM [8] uses causal intervention through a dual-branch model to mitigate clothing bias. The second category injects auxiliary ID cues, such as pose, gait, or human parsing, to strengthen discriminative features. For example, GI-ReID [9] leverages gait information to learn clothing-agnostic representations, while SCNet [10] imposes semantic consistency constraints using a human parsing network.
@@
Although these methods partially address clothing changes, they typically focus on minimizing negative clothing effects or relying on explicit cues such as body contours. We argue that visual representations inherently contain both negative factors (e.g., clothing styles that frequently change and can confuse the model) and positive factors (e.g., stable attributes such as hair, glasses, or backpacks that aid ID discrimination). As illustrated in Fig. 1(b), our goal is to remove negative influences while making implicit positive factors more explicit.
Recent advances in vision-language learning, notably Contrastive Language-Image Pre-training (CLIP) [11], have demonstrated strong abilities to bridge visual data and natural language, yielding context-aware representations beneficial for many downstream tasks [12, 13]. In ReID, methods like CLIP-ReID [14] and Instruct-ReID [15] show that aligning images with descriptive language captures rich semantics beyond appearance alone. Inspired by these successes, we leverage prompt learning to extract high-dimensional semantic cues from CLIP, thereby amplifying positive factors without additional data preprocessing.
In this paper, we introduce the Semantic Contextual Integration (SCI) framework for CC-ReID. SCI uses CLIP to capture rich semantic features and includes a Semantic Separation Enhancement (SSE) module isolates clothing-related negative factors at the text level while preserving key positive semantics for ReID. Specifically, SSE employs dual learnable text tokens to disentangle confounding semantics (both positive and negative) from clothing semantics, then orthogonalizes the token embeddings to filter out clothing factors. The refined text features guide visual encoding via our Semantic-Guided Interaction Module (SIM), focusing the model’s attention on ID-relevant cues beyond explicit contours. By integrating CLIP’s multimodal strengths, SCI enriches visual representations with high-dimensional semantic context and achieves superior performance on three standard CC-ReID benchmarks.
Our contributions are summarized threefold:
@@ II Related Work
II-A Cloth-Consistent Person ReID
Person ReID under consistent clothing conditions has been extensively studied [16]. These methods exploit clothing appearance, such as color and texture, to extract discriminative features. Feature representation learning in this context can be divided into three categories: global features, local features, and auxiliary information-based features. Global methods extract a single feature vector per image [17], while local methods aggregate part-based representations to address misalignment (e.g., human parsing [18] or horizontal partitioning [19]). Auxiliary-information approaches incorporate extra cues, such as semantic attributes [20] or synthesized samples [21], to enrich context. Strong baselines in this domain include AGW [22] and TransReID [23]. ISP [24] further refines alignment by locating body parts and carried items at the pixel level. Moreover, since person ReID is inherently formulated as a ranking problem, class imbalance can also have a significant impact on performance. Therefore, some works propose loss function modifications to handle data imbalance during training. For example, DRSL [25] optimizes the ReID model by enforcing positives to rank ahead of negatives based on distance, and further sorting positives by similarity score. Parameterized RV Loss [26] jointly optimizes retrieval and verification tasks by aligning loss functions with evaluation metrics, enabling automatic loss function search.
Although these techniques handle variations in pose, background, and viewpoint, they assume identical clothing across views. Consequently, their performance degrades when subjects change clothing, motivating research on CC-ReID.
II-B Cloth-Changing Person ReID
Several CC-ReID datasets have been proposed, PRCC [27], LTCC [28], COCAS+ [29], and VC-Clothes [30], to evaluate models under clothing variations. The primary challenge is to learn features that remain reliable despite clothing changes. Existing CC-ReID methods follow three main strategies:
II-B1 Auxiliary Soft-Biometric Cues
CAMC [31] integrates body-shape semantics into ID features, and FSAM [32] extracts fine-grained shape information to complement clothing-independent cues. M2Net [33] uses contour and parsing maps for appearance-robust features, PGAL [34] aligns keypoints via pose estimation, FLAG [35] is proposed to explicitly extract appearance and gait information, and can be integrated with most existing video-based ReID methods, and CVSL [36] jointly learns body-shape embeddings and appearance features.
II-B2 Feature Disentanglement
CAL [7] employs adversarial loss to suppress clothing-related features, and AIM [8] uses causal intervention to remove clothing bias. DCR-ReID [37] disentangles and reconstructs feature components, LDF [38] separates ID, clothing, and unrelated factors via GANs, 3DInvarReID [39] disentangles features and reconstructs two-layer 3D body shapes, FIRe2 [40] augments images with fine-grained attributes, and MAL-F [41] learns invariant features from RGB, grayscale, and contour inputs using a ResTNet backbone.
II-B3 Data Augmentation and Adaptation
Pos-Neg [42] and CCFA [43] augment clothing color and texture diversity; RCSANet [44] constructs explicit clothing-status embeddings to bolster feature robustness; MCSC [45] applies meta-learning to address clothing-distribution shifts; Zhao et al. [46] model clothing changes as fine-grained domain shifts via graph relations; and DCLR [47] synthesizes multi-clothing images via diffusion and merges them into training data.
Despite these advances, most CC-ReID methods rely solely on visual inputs, limiting their semantic depth. In this work, we integrate visual and textual modalities via CLIP to enrich semantic understanding and improve cloth-changing ReID.
@@ II-C Vision-Language Learning
Vision-language pre-training (VLP) trains models to align images with text, improving downstream visual tasks. CLIP [11] uses paired image and text encoders to learn a shared embedding space via contrastive learning, benefiting applications such as captioning and classification.
Prompt learning extends CLIP by making context tokens learnable. CoOp [48] transforms prompt words into trainable vectors, CoCoOp [49] generates input-conditional tokens per image, and DenseCLIP [50] applies pixel-text matching for dense prediction.
In person ReID, CLIP-ReID [14] introduces ID-specific tokens and a two-stage training scheme. CCLNet [51] adapts prompt learning for unsupervised visible-infrared ReID. CSDN [52] employs bimodal descriptions to align visible and infrared features. RGANet [53] uses CLIP to locate informative body parts for occluded ReID. VGSG [54] groups text features semantically to address fine-grained misalignment. CCAFL [55] integrates CLIP-generated textual information into a novel semi-supervised framework to actively learn camera-agnostic pedestrian features. MIPL [56] guides CC-ReID with multiple common prompts, and CCAF [57] leverages CLIP’s knowledge to learn fine-grained, clothing-independent semantics.
Although CLIP is not specifically pre-trained for ReID, its large-scale image-text training captures high-level semantics (e.g., body shape and context) remain invariant to clothing. We build on this by proposing CLIP-CCReID, which uses prompt learning to mitigate clothing changes without auxiliary models.
@@ III-A Preliminaries and Overview
The large-scale vision-language model CLIP [11] learns joint image-text representations via a contrastive objective. It comprises a visual encoder and a text encoder , whose embedding spaces are aligned during pre-training.
For downstream classification, CoOp [48] refines CLIP’s zero-shot prompts by introducing learnable context tokens. Specifically, given a template such as []1 []2 []M [CLS] where each []i () is a trainable vector and [CLS] is a class token, CLIP computes image-text similarities in the shared embedding space and predicts the class with maximal similarity.
Building on these ideas, we leverage CLIP and CoOp to extract rich semantic prompts for CC-ReID (see Fig. 2), enabling robust ID modeling under clothing changes.
@@ III-B Semantic Separation Enhancement
CLIP cannot explicitly ignore unwanted semantics (e.g., clothing) via negative prompts. To address this, we propose the Semantic Separation Enhancement (SSE) module, which isolates and removes clothing-related components from the high-dimensional text features, preserving only ID-relevant semantics. The overall procedure is detailed in Algorithm 1.
We define two learnable prompts:
@@
We then perform orthogonalization by subtracting the projection of the clothing feature onto the . This operation removes the clothing component that is aligned with the ID direction. This process ensures that the features are aligned with the positive features while removing the influence of negative factors. The orthogonalization process is:
where is after removing the part that is related to clothing, ensuring the feature focuses on positive information while reducing interference from clothing.
@@ III-C Semantic-Guide Interaction Module
Previous work [7, 14] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios. However, these methods typically treat visual and textual branches independently, lacking interaction. Our method refines visual representations by leveraging clothing-irrelevant textual features, enhancing the interaction between visual and textual branches. This integration introduces more robust and invariant descriptors into the visual processing, improving the model’s ability to identify individuals across clothing changes. The algorithm is summarized in Algorithm 2.
As illustrated in Fig. 3, we employ a Transformer decoder [58] to model interactions between visual and textual data. To better capture long-range dependencies within the visual domain, we replace the conventional self-attention mechanism with a non-local operation [59], which allows the model to aggregate information from spatially distant but semantically relevant regions. This capability is particularly important for handling variations in clothing. During this stage, we update the visual encoder while keeping other components frozen. The computation is as follows:
@@
To refine visual representations with semantic guidance, we leverage textual embeddings to generate attention weights over visual features. This allows the model to selectively emphasize semantically relevant regions in the visual space, leading to more discriminative and context-aware visual representations:
where is the dimensionality of the keys, is used to normalize the input features to stabilize training, and denotes a two-layer feed-forward network. This integrates visual features under the guidance of textual embeddings, and enables the model to extract the most informative visual cues aligned with the semantic content of the text.
@@
This module adopts an interaction-to-alignment approach, where text embeddings guide the visual encoding process. By emphasizing positive cues and mitigating negative ones, this method improves the model’s ability to consistently recognize individuals across different clothing scenarios.
@@
where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , is a temperature parameter, and is the set of clothes classes with the same ID. The objective of is to encourage the backbone to extract clothes-irrelevant features by penalizing its predictive power regarding clothing.
@@
By jointly optimizing , , and , the model captures ID features from easy samples (same clothes) and progressively learns to distinguish hard cases (same ID, different clothes) while aligning visual and textual representations. The weights of loss terms are basic terms [7, 14] and set as 1.
@@ IV-A Datasets and Evaluation Metrics
We evaluate our method on three standard cloth-changing datasets: LTCC [28], PRCC [27], and VC-Clothes [30], following the protocols of CAL [7] and AIM [8]. Additionally, to validate generalization, we test on two conventional ReID datasets: Market1501 [60] and MSMT17 [61].
LTCC contains 17,138 images of 152 IDs captured by 12 cameras. The training set comprises 9,576 images of 77 IDs; the testing set includes 7,543 images (493 queries and 7,050 gallery) of 75 IDs. This long-term dataset features frequent clothing changes and varied environmental conditions, with each ID wearing between two and fourteen outfits.
PRCC consists of 33,698 images of 221 IDs from three cameras. Its training set has 22,898 images of 150 IDs, and the testing set has 10,800 images of 71 IDs. Each person appears in two outfits: cameras A and B share the same clothing, while camera C uses a different outfit.
VC-Clothes is a synthetic dataset generated via the GTA5 engine. It contains 9,449 images of 256 IDs across four camera views, with 1,241 distinct clothing items distributed between training and testing sets.
Market1501 includes 12,936 training images of 751 IDs and 19,732 testing images of 750 IDs, captured by six cameras under the single-query evaluation protocol.
MSMT17 comprises 32,621 training images of 1,041 IDs and 51,027 testing images of 11,659 IDs across 15 cameras, it is one of the largest and most challenging ReID benchmarks.
We employ the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics. Rank- in CMC measures the probability of the correct match appearing among the top- retrieved images, while mAP reflects the overall retrieval performance across all queries.
Experiments are conducted under three protocols: 1) General setting: exclude same-ID and same-camera samples, using both cloth-changing and clothing-consistent pairs; 2) Same-clothes setting: exclude same-ID and same-camera samples, using only clothing-consistent pairs; 3) Cloth-changing setting: exclude same-ID, same-camera and same-clothing samples, using only cloth-changing pairs.